OpenStudy (meehan98):

Need Help Solving, Please! 8(x+1)=5(x-2)+1

OpenStudy (meehan98):

I have this so far: 8x+8=5x-10+1 8x+8=5x-9 8x=5x-17 3x=-17

OpenStudy (anonymous):

u got it right so far, so now divide 3 on both sides

OpenStudy (meehan98):

x=-5.67 or -17/3

OpenStudy (anonymous):

both are correct. depends on how you want to put it

OpenStudy (meehan98):

OKay, I just now checked it; it didn't seem right for some reason. Thanks!

OpenStudy (anonymous):

welcome :D